OpenStudy (anonymous):

Look at the formula for the area of a triangle. A = 1/2 bh Which is a solution for h?

OpenStudy (anonymous):

\[ h = 2A/b \]

OpenStudy (anonymous):

A = (1/2)bh h = A/((1/2)b) h = 2A/b

OpenStudy (anonymous):

(2A) / b = h

OpenStudy (anonymous):